# PriceFM: Foundation Model for Probabilistic Electricity Price Forecasting

### v1

@@ Conclusion
Country Region Code Price Load Solar Wind (Onshore) Wind (Offshore)
@@
Baselines
We compare the PriceFM with several spatial and temporal models. The PriceFM is optimized based on validation loss, and the hyperparameter search space is summarized in Table 2. The number of parameters of PriceFM is only 3.38M.
The spatial models, such as GraphConv, GraphAttn, GraphDiff, GraphSAGE, and GraphARMA, lack an explicit spatial decay mechanism guided by graph distance. GraphConv assigns uniform weights to all neighbors, ignoring spatial relevance; GraphAttn learns attention weights entirely from data without structural priors; GraphDiff uses diffusion kernels that spread information globally, but the importance of nodes is still determined through learned weights, without decay constraints; GraphSAGE aggregates features from sampled neighbors but lacks a notion of spatial proximity; and GraphARMA applies recursive smoothing, which can propagate noise from irrelevant neighbors. As a result, these models require large datasets to recognize spatial patterns and may struggle to suppress the influence of noisy regions. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 3. All spatial models require an adjacency matrix as input, detailed in Section Spatial Topology and Adjacency Matrix. Notably, these optimized spatial models contain more than 10M parameters.
FEDFormer, TimesNet, and TimeXer have demonstrated strong performance in general time-series forecasting tasks. FEDFormer and TimeXer are Transformer-based architectures, while TimesNet is CNN-based. However, these models are not designed for spatial graph forecasting tasks and may require large-scale data to implicitly learn spatial dependencies. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 4. Notably, the number of parameters of these optimized temporal models ranges from 5M to 10M, and they are limited to single-region prediction. As a result, 38 separate models must be trained for each output region, since they are not designed to support multi-region forecasting.
@@
Table 4: Hyperparameter search space for temporal models.
@@
Spatial Topology and Adjacency Matrix
We model the European market as a graph $G=(\mathcal{R},\mathcal{E})$, where each node $r\in\mathcal{R}$ is a bidding zone and edges indicate direct power flow via cross-border interconnections. This spatial topology is detailed in Table 5.
Let $\mathcal{N}(r)$ denote the set of directly connected neighbors of $r$, excluding $r$ itself. The binary adjacency matrix $A\in\{0,1\}^{|\mathcal{R}|\times|\mathcal{R}|}$ is defined by
@@
Metrics
We evaluate the model using Q10, Q50, Q90, AQL, AQCR, RMSE, MAE, and $R^{2}$, all averaged over regions and samples.
Q10, Q50, Q90
We compute quantile loss separately for each target quantile. For example, the Q10 loss is:
$\mathrm{Q10}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}L_{0.1}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,0.1}\right),$ (27)
and similarly for Q50 and Q90 with $\tau=0.5$ and $0.9$.
AQL
@@
AQCR
AQCR captures the proportion of forecasted distributions that violate quantile monotonicity, i.e., when a higher quantile is predicted to be smaller than a lower one. The quantile crossing indicator for a quantile pair $(\tau_{l},\tau_{u})$ with $\tau_{l}<\tau_{u}$ is defined as:
$C_{\tau_{l},\tau_{u}}(i,r,t)=\mathbb{1}\left[\hat{y}_{i,r,t,\tau_{l}}>\hat{y}_{i,r,t,\tau_{u}}\right],$ (28)
where $\mathbb{1}[\cdot]$ is the indicator function. Then, the AQCR is given by:
$\mathrm{AQCR}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}\mathrm{Cross}_{i,r,t}.$ (29)
RMSE
@@
$R^{2}$
We compute $R^{2}$ for each region and average across all regions:
$R^{2}_{r}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\bar{y}_{r}\right)^{2}},$ (34)
@@
$R^{2}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}R^{2}_{r}.$ (36)
DM Test
@@
where $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}\cdot|\mathcal{Q}|$ for probabilistic forecasts, and $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}$ for point forecasts. The index $j$ enumerates all prediction instances across dimensions, and $\hat{\sigma}_{d}$ is the sample standard deviation of $\{d_{j}\}_{j=1}^{M}$.
We compute a two-sided $p$-value using the Student’s $t$-distribution with M1M-1 degrees of freedom. A significantly positive (negative) DM value indicates that model 2 (model 1) significantly outperforms the other.

### v2

@@ Appendix C Lookup Table and Feature Availability
Country Region Code Price Load Solar
Wind
Wind
@@
Appendix D Baselines and Hyperparameters
We compare the PriceFM with several spatial and temporal models. The PriceFM is optimized based on validation loss, and the hyperparameter search space is summarized in Table 2. The number of parameters of PriceFM is only 3.38M. We use the Adam optimizer with an initial learning rate of $4\times 10^{-3}$, which decays exponentially by a factor of 0.95 every 10 epochs. The batch size is 8, which introduces a slight regularization effect. The model is trained for 50 epochs, and the checkpoint with the lowest validation loss is saved.
The graph models, such as GCN, GAT, GraphDiff, GraphSAGE, and GraphARMA, lack an explicit spatial decay mechanism guided by graph distance. GraphConv assigns uniform weights to all neighbors, ignoring spatial relevance; GraphAttn learns attention weights entirely from data without structural priors; GraphDiff uses diffusion kernels that spread information globally, but the importance of nodes is still determined through learned weights, without decay constraints; GraphSAGE aggregates features from sampled neighbors but lacks a notion of spatial proximity; and GraphARMA applies recursive smoothing, which can propagate noise from irrelevant neighbors. As a result, these models require large datasets to recognize spatial patterns and may struggle to suppress the influence of noisy regions. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 3. All spatial models require an adjacency matrix as input, detailed in Appendix E. Notably, these optimized spatial models contain more than 10M parameters.
FEDFormer, iTransformer, PatchTST, TimesNet, and TimeXer have demonstrated strong performance in general time-series forecasting tasks. FEDFormer, iTransformer, PatchTST, and TimeXer are Transformer-based architectures, while TimesNet is CNN-based. However, these models are not designed for graph forecasting tasks and may require large-scale data to implicitly learn spatial dependencies. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 4. Notably, the number of parameters of these optimized temporal models ranges from 4.96M to 10.31M, and they are limited to single-region prediction. As a result, 38 separate models must be trained for each output region, since they are not designed to support multi-region forecasting.
@@ Appendix D Baselines and Hyperparameters
Table 4: Hyperparameter search space for time-series foundation models.
@@
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
iTransformer
hidden_size: {32, 128, 512}
e_layers: {2, 3, 4}
d_ff: {512, 1024, 2048}
n_heads: {2, 4, 8}
dropout: {0.1, 0.3, 0.5}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
PatchTST
hidden_size: {32, 128, 512}
e_layers: {2, 3, 4}
n_heads: {2, 4, 8}
dropout: {0.1, 0.3, 0.5}
patch_len: {4, 6, 12}
@@
Appendix E Adjacency Matrix
We model the European market as a graph $G=(\mathcal{R},\mathcal{E})$, where each node $r\in\mathcal{R}$ is a bidding zone and edges indicate direct power flow via cross-border interconnections. This spatial topology is detailed in Table 6. Let $\mathcal{N}(r)$ denote the set of directly connected neighbors of $r$, excluding $r$ itself. The binary adjacency matrix $A\in\{0,1\}^{|\mathcal{R}|\times|\mathcal{R}|}$ is defined by
@@ Appendix F Data Scaling
Appendix G Metrics
G.1 Quantile Loss at Individual Levels
We compute quantile loss separately for each target quantile:
$\text{Q}_{\tau}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}L_{\tau}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,\tau}\right),$ (24)
where $\tau\in\{0.1,0.5,0.9\}$.
G.2 Average Quantile Loss (AQL)
@@
G.3 Average Quantile Crossing Rate (AQCR)
AQCR captures the proportion of forecasted distributions that violate quantile monotonicity, i.e., when a lower quantile is predicted to be greater than a higher one. For each sample, the quantile crossing indicator is defined as:
$C_{i,r,t}=\mathbb{I}\left(\max_{\tau_{l}<\tau_{u}}\left(\hat{y}_{i,r,t,\tau_{l}}-\hat{y}_{i,r,t,\tau_{u}}\right)>0\right)$ (25)
where $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if any quantile pair fulfills the condition inside and 0 otherwise.
We compute the AQCR as:
$\mathrm{AQCR}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}C_{i,r,t}.$ (26)
A lower AQCR indicates fewer quantile crossing violations and thus reflects more reliable probabilistic forecasts.
G.4 Root Mean Squared Error (RMSE)
@@ G.5 Mean Absolute Error (MAE)
G.6 Coefficient of Determination
We compute the Coefficient of Determination ($\text{R}^{2}$) for each region and average across all regions:
$\text{R}^{2}_{r}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\bar{y}_{r}\right)^{2}},$ (31)
@@ G.6 Coefficient of Determination
$\text{R}^{2}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\text{R}^{2}_{r}.$ (33)
G.7 Diebold & Mariano (DM) Test
@@
where $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}\cdot|\mathcal{Q}|$ for probabilistic forecasts, and $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}$ for point forecasts. The index $j$ enumerates all prediction instances across dimensions, and $\hat{\sigma}_{d}$ is the sample standard deviation of $\{d_{j}\}_{j=1}^{M}$. We compute a $p$-value; if $p<0.05$ and the DM value is positive (negative), then we report that model 2 (model 1) significantly outperforms the other in Section Experiment. The rules are summarized in Table 5.
Table 5: Interpretation of DM test outcomes.
Condition Interpretation Conclusion
$p<0.05$, DM >0>0 Statistically significant Model 2 is better
$p<0.05$, DM <0<0 Statistically significant Model 1 is better
$p\geq 0.05$ Not statistically significant